OpenStudy (anonymous):

Use Long Division to calculate: x squared minus 2x minus 30 divided by x + 7

OpenStudy (anonymous):

these are a drag

OpenStudy (anonymous):

\[\frac{x^2-2x+30}{x+7}=x-9+\frac{93}{x+7}\]

OpenStudy (anonymous):

how would you start this out

OpenStudy (anonymous):

___________________ x + 7 | x^2 - 2x + 30

OpenStudy (anonymous):

x goes in to x^2 x times then ___X________________ x + 7 | x^2 - 2x + 30 x^2 + 7x

OpenStudy (anonymous):

subtract and get ___X________________ x + 7 | x^2 - 2x + 30 x^2 + 7x -9x + 30

OpenStudy (anonymous):

x goes in to -9x -9 times ___X__- 9_____________ x + 7 | x^2 - 2x + 30 x^2 + 7x -9x + 30

OpenStudy (anonymous):

just like regular long division

OpenStudy (anonymous):

oh okay thank you :)

OpenStudy (anonymous):

yw